OpenStudy (anonymous):

sinx/(sinx+cosx)=tanx/1+tanx)

OpenStudy (anonymous):

rewrite the right hand side replacing tan with sin over cos, then simplify

OpenStudy (anonymous):

\[\frac{ \sin x }{ \sin x + \cos x } = \frac{ \tan x }{ 1+ tanx }\] L.H.S divide the whole term with cosx = \[\frac{ \frac{ \sin x }{ \cos x } }{ \frac{ sinx + cosx }{ \cos x } }\] = \[\frac{ tanx }{ tanx +1 }\] shown...

OpenStudy (anonymous):

it is aproof , right?

OpenStudy (anonymous):

Yes it is a proof!